OpenStudy (anonymous):

what is the solution to r/2=1/3? A. 2/3 B. 3/2 C. 1/6 D. 6/1

OpenStudy (zzr0ck3r):

multiply both sides by 2 and what do you have?

OpenStudy (texaschic101):

r/2 = 1/3 (cross multiply) 3r = 2 r = 2/3 or multiply both sides by 2.. 2(r/2) = 2(1/3) r = 2/3